# Numerical methods for solving PIDEs arising in swing option pricing under a two-factor mean-reverting model with jumps

### v1

@@ 7 Conclusions
In this paper, we investigated the numerical valuation of swing options with discrete action times under an affine two-factor mean-reverting model with jumps. For the numerical solution of the pertinent sequence of two-dimensional PIDEs, we studied three methods: SLCNFI, CNFI and DIRKFI. The first method is based on the semi-Lagrangian approach, whereas the second and third methods are based on a full semidiscretisation approach by suitable finite differences. For the time-stepping, the Crank-Nicolson scheme and an $L$-stable DIRK scheme are considered. Here, the nonlocal integral has been effectively treated by a fixed-point iteration.
Ample numerical experiments demonstrate that all three methods possess a desirable second-order convergence behaviour notwithstanding the convection-dominated property of the PIDE and the nonsmooth initial function and the presence of the nonlocal integral term. Our theoretical analysis confirms that the CNFI and DIRKFI methods are stable and second-order convergent under a smoothness condition. A theoretical convergence analysis of the semi-Lagrangian method is not undertaken in this work and forms an interesting topic for future research.
A main topic for our future research is the extension of the present numerical solution approach towards more advanced swing option types, such as considered in Dahlgren (2005).
@@ Acknowledgements
For the ease of presentation, write $D_{x,i,i}=g_{i}$ and assume there is index $i_{0}$ such that $g_{i_{0}}=0$. Let $P_{0}$ denote the part of the inner product $\langle D_{x}\tilde{A}_{x}V,V\rangle$ corresponding to the case $i<i_{0}$. We have by (5.1)
@@
$B$ $=-\frac{1}{2}w_{-1}\sum_{i=0}^{i_{0}-1}g_{i}\Big((V_{i-1}-V_{i})^{2}-(V_{i+1}-V_{i})^{2}\Big)-\frac{1}{2}\underbrace{(w_{-1}+w_{1})}_{=-4w_{2}}\sum_{i=0}^{i_{0}-1}g_{i}(V_{i+1}-V_{i})^{2}-\frac{1}{2}w_{2}\sum_{i=0}^{i_{0}-1}g_{i}(V_{i+2}-V_{i})^{2}$
@@
$=(2w_{-1}^{+}-4w_{2})\tilde{b}\Delta x\sum_{i=2}^{i_{0}-2}V_{i}^{2}+w_{-1}^{+}\tilde{b}\Delta x(V_{0}^{2}+2V^{2}_{1}+2V_{i_{0}-1}^{2}+V_{i_{0}}^{2})-2w_{2}\tilde{b}\Delta x(V_{1}^{2}+2V_{i_{0}-1}^{2}+2V_{i_{0}}^{2}+V_{i_{0}+1}^{2})$
$\qquad-\frac{1}{2}w_{-1}g_{0}(V_{0}-V_{-1})^{2}+w_{2}g_{0}(V_{1}-V_{0})^{2}.$

### v2

@@ 7 Conclusions
In this paper, we investigated the numerical valuation of swing options with discrete action times under an affine two-factor mean-reverting model with jumps. For the numerical solution of the pertinent sequence of two-dimensional PIDEs, we studied three methods: SLCNFI, CNFI and DIRKFI. The first method is based on the semi-Lagrangian approach, whereas the second and third methods are based on a full semidiscretisation approach by suitable finite differences. For the time-stepping, the Crank–Nicolson scheme and an $L$-stable DIRK scheme are considered. Here, the nonlocal integral has been effectively treated by a fixed-point iteration.
Ample numerical experiments demonstrate that all three methods possess a desirable second-order convergence behaviour notwithstanding the convection-dominated property of the PIDE and the nonsmooth initial function and the presence of the nonlocal integral term. Our theoretical analysis confirms that the CNFI and DIRKFI methods are stable and second-order convergent under a smoothness condition. A theoretical convergence analysis of the semi-Lagrangian method is not undertaken in this work and forms an interesting topic for future research. The finite difference framework adopted for swing option valuation enables also the efficient approximation of the Delta Greeks, which provides important insight into the impact of the price components $x$ and $y$ on the option values. Moreover, a study of the effect of the financial model parameters on the optimal exercise policy sheds much light on when and how much to exercise.
Among others, a main topic for our future research is the extension of the present numerical solution approach towards more advanced swing option types, such as considered in Dahlgren (2005).
@@ Acknowledgements
For the ease of presentation, write $D_{x,i,i}=:g_{i}$ and assume there is index $i_{0}$ such that $g_{i_{0}}=0$. Let $P_{0}$ denote the part of the inner product $\langle D_{x}\tilde{A}_{x}V,V\rangle$ corresponding to the case $i<i_{0}$. We have by (5.1)
@@
$B$ $=-\frac{1}{2}w_{-1}\sum_{i=0}^{i_{0}-1}g_{i}\Big((V_{i-1}-V_{i})^{2}-(V_{i+1}-V_{i})^{2}\Big)-\frac{1}{2}\underbrace{(w_{-1}+w_{1})}_{=-4w_{2}}\sum_{i=0}^{i_{0}-1}g_{i}(V_{i+1}-V_{i})^{2}$
$\qquad-\frac{1}{2}w_{2}\sum_{i=0}^{i_{0}-1}g_{i}(V_{i+2}-V_{i})^{2}$
@@
$=(2w_{-1}^{+}-4w_{2})\tilde{b}\Delta x\sum_{i=2}^{i_{0}-2}V_{i}^{2}+w_{-1}^{+}\tilde{b}\Delta x(V_{0}^{2}+2V^{2}_{1}+2V_{i_{0}-1}^{2}+V_{i_{0}}^{2})$
$\qquad-2w_{2}\tilde{b}\Delta x(V_{1}^{2}+2V_{i_{0}-1}^{2}+2V_{i_{0}}^{2}+V_{i_{0}+1}^{2})-\frac{1}{2}w_{-1}g_{0}(V_{0}-V_{-1})^{2}+w_{2}g_{0}(V_{1}-V_{0})^{2}.$